### v1

# Highly Efficient Self-Adaptive Reward Shaping for Reinforcement Learning

Abstract
Reward shaping addresses the challenge of sparse rewards in reinforcement learning by constructing denser and more informative reward signals. To achieve self-adaptive and highly efficient reward shaping, we propose a novel method that incorporates success rates derived from historical experiences into shaped rewards. Our approach utilizes success rates sampled from Beta distributions, which dynamically evolve from uncertain to reliable values as more data is collected. Initially, the self-adaptive success rates exhibit more randomness to encourage exploration. Over time, they become more certain to enhance exploitation, thus achieving a better balance between exploration and exploitation. We employ Kernel Density Estimation (KDE) combined with Random Fourier Features (RFF) to derive the Beta distributions, resulting in a computationally efficient implementation in high-dimensional continuous state spaces. This method provides a non-parametric and learning-free approach. The proposed method is evaluated on a wide range of continuous control tasks with sparse and delayed rewards, demonstrating significant improvements in sample efficiency and convergence stability compared to several baselines.
1 Introduction
Environments with sparse and delayed rewards present notable challenges for reinforcement learning (RL), despite its significant progress in various fields. In such contexts, as the reward structure lacks immediate signals, agents receive feedback only after long horizons, making the ability to quickly and effectively discover beneficial samples crucial for successful learning (Ladosz et al., 2022). To address this, a straightforward solution is to reconstruct the reward models by introducing auxiliary signals to assess the agent’s behavior at each step, which has led to the popular technique of Reward Shaping (RS) (Sorg et al., 2010a, b; Gupta et al., 2022). Inverse reinforcement learning (IRL), which extracts reward functions from human knowledge or expert demonstrations, represents an intuitive approach within this framework. However, this heavily relies on extensive human input and data, which can be difficult to obtain especially in complex environments. Alternatively, fully autonomous approaches have emerged as an attractive direction.
Automatically maintained shaped rewards can be generally categorized into two branches: intrinsic motivation based signals, which are task-agnostic, and hidden value based signals, which are commonly task-related. The former mainly introduces exploration bonuses to encourage agents to explore a wider range of states, typically by rewarding novel or less frequently visited states (Ostrovski et al., 2017; Devidze et al., 2022; Tang et al., 2017; Burda et al., 2018; Bellemare et al., 2016). These methods have shown effective exploration enhancement, however, often overlooking the internal values of the states themselves. This can lead to the "noisy TV" problem, where agents become fixated on highly novel but meaningless regions, thus trapping agents in suboptimal behaviors (Mavor-Parker et al., 2022). In contrast, the latter leverages high-level heuristics to guide agents in retrieving hidden values from their experiences, which can significantly improve the agent’s convergence (Trott et al., 2019; Ma et al., 2024; Zheng et al., 2018).
To overcome the limitations of existing RS methods, and try to combine the advantages of exploration-encouraged and hidden value-based rewards, this paper proposes a Self-Adaptive Success Rate based reward shaping mechanism (SASR) [^1] The source code is accessible at: https://github.com/mahaozhe/SASR. The success rate, defined as the ratio of a state’s presence in successful trajectories to its total occurrences, works as an auxiliary reward distilled from historical experience. The success rate assesses a state’s contribution toward successful task completion, which aligns with the agent’s original objectives, offering informative guidance for agent learning. Furthermore, to mitigate overconfidence caused by deterministic success rates, we adopt Beta distributions to model success rates from a probabilistic perspective. Beta distributions allow for a self-adaptive evolution of confidence in approximating a state’s success rate. This effectively prevents premature certainty and ensures stable convergence to reliable rewards as more data accumulates. To derive Beta distributions, we leverage kernel density estimation enhanced by random Fourier features to efficiently estimate success and failure counts, formulating a highly efficient approach. The main contributions of this paper are summarized as follows:
  1. (i)
    We propose SASR, an autonomous reward-shaping mechanism without the requirement of expert demonstrations or domain knowledge. By extracting a metric from historical experiences that aligns with the agent’s optimization objectives, SASR effectively densifies the environmental rewards.
  2. (ii)
    We introduce a novel self-adaptive strategy to achieve a balance between exploration and exploitation. Initially, low-confidence Beta distributions provide stochastic rewards that encourage exploration by diversifying the policy optimization directions. With the accumulation of experience, high-confidence Beta distributions offer more reliable and precise rewards to enhance exploitation.
  3. (iii)
    To derive the Beta distributions in continuous state spaces, we implement kernel density estimation, leveraging random Fourier features. This formulates a non-parametric and learning-free approach, thereby achieving remarkably low computational complexity.
  4. (iv)
    We use sparse-reward and continuous control environments, MuJoCo (Todorov et al., 2012) and simulated robot (de Lazcano et al., 2023), to evaluate the performance of SASR. It significantly outperforms several baselines in terms of sample efficiency, learning speed, and convergence stability.
2 Related Work
Reward shaping (RS) techniques can be broadly categorized based on the source of learning: either from human knowledge or agent’s own experiences. Techniques drawing reward models from human knowledge rely on pre-existing expertise, where methods such as Inverse Reinforcement Learning (IRL) (Arora and Doshi, 2021; Ramachandran and Amir, 2007; Ziebart et al., 2008; Hadfield-Menell et al., 2016) and Inverse Optimal Control (IOC) (Schultheis et al., 2021) aim to derive reward or objective functions from expert demonstrations. Another branch within this approach, Imitation Learning (IL) (Ho and Ermon, 2016; Duan et al., 2017; Brown et al., 2020), attempts to clone human behaviors from offline data without explicitly modeling reward structures. Following this, transferring the extracted reward models to new tasks has received considerable efforts (Bıyık et al., 2022; Wu et al., 2021; Ellis et al., 2021; Cheng et al., 2021; Adamczyk et al., 2023). However, these methods require significant human-driven data and generally show limited adaptability to out-of-distribution (OOD) scenarios. Therefore, we focus instead on the autonomous self-learning approaches, which based on the meaning of rewards, can be discussed from two perspectives: intrinsic motivation based and hidden value based rewards.
Intrinsic motivation based RS explores general heuristics or task-agnostic metrics to encourage exploration in sparse-reward environments. Potential-based algorithms define the shaped reward in the form of $\gamma\Phi(s^{\prime})-\Phi(s)$, where $\Phi(\cdot)$ is a potential function, enabling the reward to be nullified in the Bellman equation, thus ensuring the optimal policy remains invariant (Devlin and Kudenko, 2012; Asmuth et al., 2008). However, the potential function’s design heavily depends on environmental dynamics, making it more suitable in model-based RL. More popular methods incorporate exploration bonuses by rewarding novel states (Devidze et al., 2022; Hong et al., 2018; Eysenbach et al., 2019). For instance, count-based strategies maintain visitation counts and assign higher rewards to less frequently visited states. To overcome the challenge of state counting in continuous spaces, Tang et al. (2017) utilized a hash function to discretize the state space, Bellemare et al. (2016) introduced pseudo-counts based on recording probabilities, and Ostrovski et al. (2017) used PixelCNN Van den Oord et al. (2016). Our work adopts a KDE approach to estimate the state counts. Additionally, Burda et al. (2018) proposed random network distillation to assess state novelty in high-dimensional spaces; Curiosity-driven approaches reward encounters with states that are surprising and out-of-prediction (Sun et al., 2022; Pathak et al., 2017). The exploration bonuses branch shows effective exploration enhancement, but only considering novelty while ignoring the underlying values may lead to sub-optimal behaviors. A notable example illustrating this limitation is the "noisy TV" problem (Mavor-Parker et al., 2022), where agents may become fixated on irrelevant and dynamic stimuli that yield high novelty but no useful information.
Hidden value based RS, on the other hand, primarily focuses on task-related signals to reveal the extent to which states contribute to achieving higher environmental rewards and their underlying significance. Trott et al. (2019) integrated the distance between a state and the target as additional rewards. Stadie et al. (2020) derived informative reward structures using a Self-Tuning Network (STN) to optimize reward guidance. Memarian et al. (2021) captured the preferences among different trajectories by ranking them via a trained classifier. Zheng et al. (2018) minimized the KL-divergence between the learned and the original environmental rewards to synchronize their distributions. Mguni et al. (2023) used an auxiliary agent that competed against the original agent in a Markov game; Instead, Ma et al. (2024) introduced ReLara, a collaborative framework where an assistant reward agent automatically generates rewards to guide the policy agent. Moreover, the incorporation of multiple agents or hierarchical structures to share and transfer knowledge through synchronized reward functions is also a popular research direction (Park et al., 2023; Yi et al., 2022; Gupta et al., 2023; Hu et al., 2020).
While intrinsic motivation rewards directly encourage exploration, hidden value rewards support efficient convergence, Ma et al. (2024) combined these two advances by introducing a scheme that evolves the shaped rewards from random perturbations to meaningful metrics. Injecting random rewards during early training relaxes the assumption that only novel states should be rewarded, suggesting instead that any state can be randomly rewarded. Building on this concept, our study considers the success rate, which is the probability that a state will contribute to task completion, as the hidden value. We sample this metric from evolving Beta distributions that progressively increase their confidence, thus aligning the process to balance the exploration and exploitation adaptively. Compared with ReLara, which relies on a black-box agent, our method derives the shaped rewards directly from historical data, effectively reducing the delay of reward convergence.
3 Preliminaries
Reinforcement Learning (RL) aims to train an agent to interact with an environment, which is commonly modeled as a Markov Decision Process (MDP). An MDP represented as $\langle S,A,T,R,\gamma\rangle$, involves four main components: $S$ is the state space, $A$ is the action space, $T:S\times A\times S\rightarrow[0,1]$ is the probability of transitioning from one state to another given a specific action, and $R:S\rightarrow\mathbb{R}$ is the reward model. The objective in RL is to learn a policy $\pi(a|s)$ that maximizes the expected cumulative rewards $G=\mathbb{E}[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t})]$, where $\pi(a|s)$ indicates the likelihood of selecting action $a$ in state $s$, and $\gamma$ is the discount factor. Model-free RL methods include updating value functions, directly optimizing policies, or employing an actor-critic framework (Sutton and Barto, 2018).
Beta Distribution is defined on the interval $[0,1]$, making it ideal for modeling proportions or probabilities. It is parameterized by $\alpha$ and $\beta$, which are interpreted as prior counts of successes and failures of a binary outcome. The probability density function of a Beta-distributed variable $X$ is:
$f(x;\alpha,\beta)=\frac{1}{B(\alpha,\beta)}x^{\alpha-1}(1-x)^{\beta-1},$ (1)
where $B(\alpha,\beta)$ is a normalization factor. The key characteristic of Beta distribution is its adaptability: as more data accumulates, the values of $\alpha$ and $\beta$ increase, the distribution’s shape becomes narrower, thereby increasing confidence in the estimated probabilities. This feature is particularly beneficial in adaptive online learning, aligning closely with our objective of balancing exploration and exploitation.
Kernel Density Estimation (KDE) is a non-parametric approach to approximate the probability density function of a random variable from data samples. Given $n$ data points $\{x_{i}\}_{i=1}^{n}$, KDE smooths these points to approximate the density function as follows:
$\hat{d}(x)=\frac{1}{nh}\sum_{i=1}^{n}K\left(\frac{x-x_{i}}{h}\right),$ (2)
where $h$ is the bandwidth, $K(\cdot)$ is a kernel function such as Gaussian kernel, Laplacian kernel, or Cauchy kernel. KDE is particularly useful in scenarios where the actual distribution is not well-defined or complex, such as continuous state spaces in our RL environments.
4 Methodology
We propose a Self-Adaptive Success Rate based reward shaping mechanism (SASR) to accelerate RL algorithms in sparse-reward environments. Figure 1 illustrates the principles of the SASR mechanism: The diagram consists of two parts representing the early and late stages of learning. From the early stage to the late stage, the Beta distributions modeling the estimation of the true success rates evolve from being more stochastic to more deterministic as experiences accumulate. This progression closely aligns with the agent’s exploration-exploitation balance. The process of Beta distributions’ evolution and generation of shaped rewards from them are detailed in Section 4.1. Additionally, to achieve highly efficient computation, we leverage KDE and RFF to estimate success and failure counts, which are used to derive the corresponding Beta distributions, as detailed in Section 4.2. Lastly, Section 4.3 presents the integration of SASR into the RL agent and the overall algorithmic flow.
Figure 1: A schematic diagram of the self-adaptive success rate based reward shaping mechanism. KDE: Kernel Density Estimation; RFF: Random Fourier Features.
4.1 Self-Adaptive Success Rate Sampling
We consider the following augmented reward function for SASR:
$R^{SASR}(s)=R^{E}(s)+\lambda R^{S}(s),$ (3)
where $R^{E}(s)$ is the environmental reward, $R^{S}(s)$ is the additional shaped reward, and $\lambda$ is a weight factor to control its relative scale. When $\lambda\rightarrow 0$, we can recover the original reward model. We assign the shaped reward $R^{S}(s_{i})$ of a given state based on its success rate – defined as the ratio of the state’s presence in successful trajectories to its total occurrences. This metric provides a meaningful and informative reward from a statistical perspective, as a higher success rate indicates a greater likelihood and contribution for the state to guide the agent towards successful task completion, thus deserving a higher shaped reward. Formally, the success rate based shaped reward $R^{S}(s_{i})$ is given by:
$R^{S}(s_{i})=f\left(\frac{N_{S}(s_{i})}{N_{S}(s_{i})+N_{F}(s_{i})}\right),$ (4)
where $N_{S}(s_{i})$ and $N_{F}(s_{i})$ denote the counts of appearances for state $s_{i}$ in historical successful and failed trajectories respectively, and $f(\cdot)$ is a mapping function that adjusts the magnitude of rewards: $f:[0,1]\rightarrow[R_{\min},R_{\max}]$.
Given the counts $N_{S}(s_{i})$ and $N_{F}(s_{i})$ derived from prior experiences, directly using a deterministic success rate $\frac{N_{S}(s_{i})}{N_{S}(s_{i})+N_{F}(s_{i})}$ may lead to overconfidence in the estimation of the true value. To address this, we model the success rate estimation from a probabilistic perspective, inspired by the principles of Thompson sampling, an effective algorithm widely used to balance exploration and exploitation in multi-armed bandit problems (Thompson, 1933; Agrawal and Goyal, 2012). We model the success rate of each state as an approximated variable in a Beta distribution, setting the shape parameters $\alpha=N_{S}(s_{i})+1$ and $\beta=N_{F}(s_{i})+1$:
$r^{S}_{i}\sim\text{Beta}(r;\alpha,\beta)=\frac{1}{B(N_{S}(s_{i})+1,N_{F}(s_{i})+1)}r^{N_{S}(s_{i})}(1-r)^{N_{F}(s_{i})},$ (5)
where the beta function $B(\cdot,\cdot)$ is the normalization factor. By sampling from this distribution, we obtain a probabilistic estimate of the state’s success rate. This sampled value, $r^{S}_{i}$, is then processed through the mapping function $f(\cdot)$ to produce the shaped reward: $R^{S}(s_{i})=f(r^{S}_{i})$.
As $N_{S}(s_{i})$ and $N_{F}(s_{i})$ progressively increase throughout the learning process, they influence the shape and sampling variability of the Beta distribution. Generating the shaped reward from these evolving Beta distributions offers several advantages:
  • Encourage Exploration. In the early stages of learning, the counts $N_{S}(s_{i})$ and $N_{F}(s_{i})$ are relatively low, resulting in broader Beta distributions and greater variability in sampled values, which injects significant randomness into the shaped rewards. These random rewards optimize the agent in various directions with tiny steps, thus shifting the anchors from which the stochastic policy samples actions. The strategy effectively broadens the range of policies and actions, thus encouraging the exploration. (Supporting experimental results are provided in Section 5.2.)
  • Enhance Exploitation. In the late stage of learning, as the counts $N_{S}(s_{i})$ and $N_{F}(s_{i})$ increase, the Beta distribution gradually sharpens, concentrating generated rewards closer to the true success rate. The evolution to a certain reward signal with higher confidence highly supports the agent’s exploitation phase, facilitating more efficient convergence towards optimal policies.
  • Consistent Optimization. The peak of the Beta distribution, $B(N_{S}(s_{i})+1,N_{F}(s_{i})+1)$, is given by $\frac{N_{S}(s_{i})}{N_{S}(s_{i})+N_{F}(s_{i})}$, directly approximating the success rate. Meanwhile, the expected value of the distribution, $\frac{N_{S}(s_{i})+1}{N_{S}(s_{i})+N_{F}(s_{i})+2}$, aligns closely with the calculated success rate. This ensures that, despite some randomness, the overall reward remains consistent with policy optimization.
4.2 Highly Efficient Beta Distribution Derivation
In this section, we present the methods to obtain the success and failure counts, $N_{S}(s_{i})$ and $N_{F}(s_{i})$, which are required for deriving the Beta distributions. In tabular environments with low-dimensional, discrete, and finite state spaces, tracking these counts can be straightforwardly accomplished using two tables. However, most real-world environments involve high-dimensional, continuous, and infinite state spaces, making it impractical to maintain exact counts for every state.
To address this challenge, we use Kernel Density Estimation (KDE) to approximate the densities of successes and failures from accumulated experience. Specifically, we maintain two buffers, $\mathcal{D}_{S}$ and $\mathcal{D}_{F}$, to store the states in successful and failed trajectories, respectively. By treating these states as scattered data instances distributed across the state space, KDE estimates the density as:
$\tilde{d}_{X}(s_{i})=\frac{1}{|\mathcal{D}_{X}|}\sum_{j=1}^{|\mathcal{D}_{X}|}{K(s_{i}-s_{j})},\quad X\in\{S,F\},$ (6)
where $K(\cdot)$ is the kernel function, and $|\mathcal{D}_{X}|$ is the total number of states in the corresponding buffer. We select the Gaussian kernel in our implementation. The estimated density $\tilde{d}_{X}(s_{i})$ serves as a proxy for the likelihood of encountering state $s_{i}$ under success or failure scenarios, providing a statistically sound basis for estimating $N_{X}(s_{i})$. By multiplying $\tilde{d}_{X}(s_{i})$ with the total number of observed states $N$, and representing states as $k$-dimensional vectors, the count $\tilde{N}_{X}(s_{i})$ is estimated as:
$\tilde{N}_{X}(\bm{s}_{i})=N\times\tilde{d}_{X}(\bm{s}_{i})=\frac{N}{|\mathcal{D}_{X}|}\sum_{j=1}^{|\mathcal{D}_{X}|}{\exp\left(-\frac{||\bm{s}_{i}-\bm{s}_{j}||^{2}}{2\sigma^{2}}\right)},\quad X\in\{S,F\},$ (7)
where $\sigma$ is a hyperparameter that controls the bandwidth of the Gaussian kernel.
We further integrate Random Fourier Features (RFF) (Rahimi and Recht, 2007) to reduce computational complexity, as calculating the Gaussian kernel can be expensive, especially in scenarios involving high-dimensional state spaces and large buffers. RFF approximates the kernel function of the original $k$-dimensional states through an inner product of $M$-dimensional randomized features:
$K(\bm{s}_{i},\bm{s}_{j})\approx z(\bm{s}_{i})^{T}z(\bm{s}_{j}),\quad z(\bm{s})=\sqrt{\frac{2}{M}}\cos(\bm{W}^{T}\bm{s}+\bm{b}),$ (8)
where $z(\cdot)$ is the RFF mapping function with $\bm{W}=\big{[}\bm{w}^{(1)},\ldots,\bm{w}^{(M)}\big{]}\in\mathbb{R}^{k\times M}$ and $\bm{b}=\big{[}b^{(1)},\ldots,b^{(M)}\big{]}^{T}\in\mathbb{R}^{M}$ randomly sampled from the following spectral distributions:
$\bm{w}^{(m)}\sim\mathcal{N}(\bm{0},\sigma^{-2}\bm{I}_{k}),\quad b^{(m)}\sim\text{Uniform}(0,2\pi),\quad m=1,\ldots,M,$ (9)
where $\bm{I}_{k}$ is the $k\times k$ identity matrix. Equation 9 is applied for the Gaussian kernel, while different kernels, such as Laplacian or Cauchy kernels, have different spectral distributions. Detailed derivations of the RFF method and the list of spectral distributions are provided in Appendix A.
Suppose for each iteration, the buffer size of $\mathcal{D}_{X}$ is $D$ and the batch size is $B$, the computational complexity to retrieve the corresponding count $N_{X}$ is $O(MDB)$, representing a linear complexity (detailed derivation in Appendix B). RFF transforms the nonlinear kernel computations into vectors’ linear operations, significantly accelerating computation by leveraging the vectorization capabilities of GPUs (Dongarra et al., 2014). This demonstrates its higher efficiency compared to other methods that rely on network updates and inferences involving extensive nonlinear computations.
Moreover, we introduce a hyperparameter, the retention rate $\phi\in(0,1]$, to regulate the volume and diversity of states stored in the buffers. Instead of storing all the states in each trajectory, we uniformly retain a specific portion $\phi$. This strategy offers two main benefits: (1) as adjacent states in one trajectory tend to be highly similar, especially those closer to the initial state are repetitive and uninformative, retaining a portion of states can skip redundant states and increase sample diversity; (2) maintaining a lower retention rate in the early training stage keeps lower $N_{S}$ and $N_{F}$, resulting in a broader Beta distribution and avoiding premature overconfidence.
Given the retention rate $\phi$, the space complexity of maintaining two buffers is $O(\phi T|s|)$, where $T$ is the total number of iterations and $|s|$ is the size of a single state. To address scalability concerns in more complex or high-dimensional scenarios, we propose leveraging the existing replay buffer of the backbone algorithm, soft actor-critic. By augmenting this buffer with a flag array that indicates whether each state is a success or failure, while this introduces additional indexing computations during processing, we can efficiently save space requirements.
4.3 The SASR Mechanism for RL agents
Algorithm 1 Self-Adaptive Success Rate based Reward Shaping
0:  Environment $\mathcal{E}$ and agent $\mathcal{A}$.
0:  Experience replay buffer $\mathcal{D}$.
0:  State buffers for success $\mathcal{D}_{S}$ and failure $\mathcal{D}_{F}$.
0:  RFF mapping function $z:\mathbb{R}^{k}\rightarrow\mathbb{R}^{M}$. $\triangleright$ Sample $\bm{W}$ and $\bm{b}$ based on Equation 9
1:  for each trajectory $\tau=\emptyset$ do
2:     for each environmental step do
3:        $(s_{t},a_{t},s_{t+1},r^{E}_{t})\leftarrow\text{CollectTransition}(\mathcal{E},\mathcal{A})$triangleright interact with the environment
4:        $\mathcal{D}\leftarrow\mathcal{D}\cup\{(s_{t},a_{t},s_{t+1},r^{E}_{t})\}$triangleright store the transition in the replay buffer
5:        $\tau\leftarrow\tau\cup\{s_{t}\}$triangleright record the state in the trajectory
6:     end for
7:     if trajectory is successful: $\mathcal{D}_{S}\leftarrow\mathcal{D}_{S}\cup\tau$triangleright store the trajectory in the success buffer
8:     else: $\mathcal{D}_{F}\leftarrow\mathcal{D}_{F}\cup\tau$triangleright otherwise, store the trajectory in the failure buffer
9:  end for
10:  for each update step do
11:     $\{(s_{t},a_{t},r^{E}_{t},s_{t+1})_{i}\}\sim\mathcal{D}$triangleright sample a batch of transitions from the replay buffer
12:     $\tilde{N}_{S}=z(\bm{s}_{t})^{T}z(\bm{\mathcal{D}_{S}})$triangleright estimate the success counts
13:     $\tilde{N}_{F}=z(\bm{s}_{t})^{T}z(\bm{\mathcal{D}_{F}})$triangleright estimate the failure counts
14:     $r^{S}_{t}\sim\text{Beta}(r;\tilde{N}_{S}+1,\tilde{N}_{F}+1)$triangleright sample the success rate from the Beta distribution
15:     $r^{SASR}_{t}=r^{E}_{t}+\lambda f(r^{S}_{t})$triangleright compute the SASR reward
16:     Update agent $\mathcal{A}$ with $\{(s_{t},a_{t},r^{SASR}_{t},s_{t+1})_{i}\}$
17:  end for
Building upon the SASR reward, we employ the soft actor-critic (SAC) algorithm of Haarnoja et al. (2018a) as the foundation for our RL agent. Let $\pi^{\mathcal{A}_{P}}_{\theta}$ be parameterized policy network and $Q^{\mathcal{A}_{P}}_{\phi}$ be parameterized Q-network. We optimize the Q-function by minimizing the mean squared error (MSE) between the predicted Q-value and the temporal difference (TD) target:
$\mathcal{L}(\phi)=\Big{(}Q_{\phi}(s_{t},a_{t})-\big{(}r^{E}_{t}+\lambda R^{S}(s_{t})+\gamma Q_{{\phi}^{\prime}}(s_{t+1},a_{t+1})\big{)}\Big{)}^{2},$ (10)
where $Q_{{\phi}^{\prime}}$ is obtained from a secondary frozen target network to maintain a fixed objective (Mnih et al., 2015). It is worth noting that the environmental reward $r^{E}_{t}$ is retrieved from the replay buffer. Conversely, the shaped reward $R^{S}(s_{t})$ is computed in real-time using the most recently updated $N_{S}(s_{t})$ and $N_{F}(s_{t})$, thereby capturing the latest advancements in learning progress.
We optimize the policy network by maximizing the expected Q-value and the entropy of the policy $\mathcal{H}\big{(}\pi^{\mathcal{A}_{P}}(\cdot|s_{t})\big{)}$ (Haarnoja et al., 2018b):
$\mathcal{L}(\theta)=\mathbb{E}_{a_{t}\sim\pi_{\theta}(\cdot|s_{t})}\big{[}-Q_{\phi}(s_{t},a_{t})+\log{\pi_{\theta}(a_{t}|s_{t})}\big{]}.$ (11)
The outline of the SASR is summarized in Algorithm 1. We maintain two buffers to store successful and failed states and estimate the success and failure counts from them. Subsequently, shaped rewards are computed by sampling from Beta distributions followed by a scale transformation. The shaped rewards, combined with environmental rewards, are used to update the agent’s Q-function and policy.
In summary, SASR offers two primary advantages: First, it leverages the adaptive properties of Beta distributions to continuously adjust the agent’s confidence levels in understanding the environment, effectively maintaining the balance between exploration and exploitation. Second, SASR operates as a non-parametric, learning-free approach, enabling highly efficient reward calculation.
5 Experiments
We evaluate SASR on continuous control environments, including seven MuJoCo tasks (Todorov et al., 2012): AntStand, AntSpeed, AntFar, AntVerFar, WalkerKeep, HumanStand, and HumanKeep, two robotic tasks (de Lazcano et al., 2023): RobotReach and RobotPush, as well as one physical simulation task MountainCar (Towers et al., 2023), shown in Figure 2. Our environment setup follows the guidelines established in the ReLara paper (Ma et al., 2024), with ReLara also included as a baseline for comparison. The tasks provide only sparse rewards, with a reward of $1$ granted upon reaching the final objective within the maximum permitted steps, and 00 for all other states. To ensure robust experimental validation, we train multiple instances for each setting using different random seeds and report the average results. Additionally, we maintain uniform hyperparameters and network architectures across all experiments, as detailed in Appendix D.
AntStand
AntSpeed
AntFar
AntVeryFar
WalkerKeep
HumanKeep
HumanStand
RobotReach
RobotPush
MountainCar
Figure 2: MuJoCo, robotic and physical simulation tasks with their objectives in our experiments.
5.1 Comparison and Discussion
Baselines. We compare SASR with eight baselines to benchmark its performance: (a) RL with an Assistant Reward Agent (ReLara) (Ma et al., 2024), (b) RL Optimizing Shaping Algorithm (ROSA) (Mguni et al., 2023), (c) Exploration-Guided RS (ExploRS) (Devidze et al., 2022), (d) Count-based static hashing exploration (#Explo) (Tang et al., 2017), (e) Random Network Distillation (RND) (Burda et al., 2018), (f) Soft Actor-Critic (SAC) (Haarnoja et al., 2018a), (g) Twin Delayed Deep Deterministic Policy Gradient (TD3) (Fujimoto et al., 2018), and (h) Proximal Policy Optimization (PPO) (Schulman et al., 2017). Algorithms (a) to (e) are all reward shaping methods, involving either exploration bonuses or auxiliary agents to provide additional rewards to tackle the sparse-reward challenge, while algorithms (f) to (h) are advanced RL algorithms. For the implementation of these baselines, we employ either the CleanRL library (Huang et al., 2022) or codes provided by the authors.
Figure 3 and Table 1 report the average episodic returns and standard errors throughout training. Our findings indicate that SASR surpasses the baselines in terms of sample efficiency, learning stability, and convergence speed. The primary challenges in these environments are attributed to delayed and sparse rewards alongside long horizons, making exploration essential for timely obtaining successful trajectories. Although the exploration strategies of algorithms such as ExploRS, #Explo, and RND are designed to reward novel states, effectively expanding the early exploration space with the direct additional target, they continue focusing on discovering novel states, overlooking the implicit values of these states, which makes them fail to return to the final objectives.
Figure 3: The learning performance of SASR compared with the baselines.
Table 1: The average episodic returns and standard errors of SASR and the baselines.
Algorithms AntStand AntSpeed AntFar AntVeryFar WalkerKeep HumanStand HumanKeep RobotReach RobotPush MountainCar
SASR 39.12$\pm$2.86 0.94$\pm$0.07 73.92$\pm$4.97 78.64$\pm$3.92 158.24$\pm$5.59 42.63$\pm$2.17 180.98$\pm$4.40 81.29$\pm$6.52 137.06$\pm$12.66 0.91$\pm$0.04
ReLara 28.66$\pm$1.82 0.33$\pm$0.02 67.77$\pm$4.30 64.07$\pm$4.17 77.14$\pm$8.77 29.72$\pm$1.85 160.31$\pm$7.30 103.56$\pm$7.18 58.71$\pm$6.98 0.89$\pm$0.01
ROSA 3.80$\pm$0.03 0.02$\pm$0.00 4.71$\pm$0.28 0.64$\pm$0.07 32.14$\pm$1.19 8.55$\pm$0.03 152.38$\pm$4.98 0.27$\pm$0.03 $0$0.90$\pm$0.02
ExploRS 4.52$\pm$0.04 0.01$\pm$0.00 5.42$\pm$0.22 1.67$\pm$0.10 2.47$\pm$0.13 8.63$\pm$0.03 158.09$\pm$4.42 0.79$\pm$0.04 0.20$\pm$0.08 -0.99$\pm$0.02
#Explo 6.79$\pm$0.50 0.00$\pm$0.00 12.40$\pm$1.66 0.55$\pm$0.09 131.56$\pm$5.40 28.73$\pm$1.79 160.60$\pm$7.04 4.19$\pm$0.42 6.31$\pm$0.85 0.79$\pm$0.02
RND 4.23$\pm$0.03 0.02$\pm$0.00 6.66$\pm$0.25 1.00$\pm$0.09 44.78$\pm$1.39 8.67$\pm$0.03 159.79$\pm$4.27 28.18$\pm$2.53 0.04$\pm$0.04 0.94$\pm$0.00
SAC 15.93$\pm$0.69 0.00$\pm$0.00 9.64$\pm$0.57 20.73$\pm$2.87 70.96$\pm$8.10 9.31$\pm$0.05 4.59$\pm$0.84 45.03$\pm$4.92 0.55$\pm$0.21 -0.05$\pm$0.02
TD3 0.00$\pm$0.00 0.00$\pm$0.00 0.81$\pm$0.02 0.81$\pm$0.02 18.62$\pm$0.75 5.72$\pm$0.04 0.55$\pm$0.03 0.00$\pm$0.00 0.00$\pm$0.00 0.00$\pm$0.00
PPO 4.54$\pm$0.04 0.00$\pm$0.00 6.33$\pm$0.24 1.80$\pm$0.11 33.77$\pm$1.11 8.36$\pm$0.03 138.13$\pm$12.64 79.52$\pm$10.80 0.00$\pm$0.00 0.93$\pm$0.00
SASR, on the other hand, utilizes the success rate as auxiliary information to assess each state’s contribution to task completion. This metric is consistent with the agent’s primary objectives, and incorporating this meaningful signal into the temporal difference updating at each step significantly enhances learning efficiency. Without this integration, the agent would have to wait until the end of the trajectory to obtain informative signals. Furthermore, SASR regulates an efficient exploration strategy by injecting substantial random rewards during the early phases to optimize the agent in multiple directions thereby increasing the likelihood of collecting positive samples. This is achieved by sampling from initially wider Beta distributions. Importantly, these Beta distributions adaptively yield increasingly confident and accurate success rates as more data is collected, thereby accelerating the agent’s learning. These strategies collectively enhance SASR’s efficiency and stability in handling these challenging tasks.
While ReLara employed a similar exploration mechanism with SASR by rewards perturbations to encourage exploration, it relies on a separate black-box agent, which requires more iterations to converge. In contrast, SASR’s success rate sampling is more intuitive and direct, reducing the delay in providing valuable information to the policy agent. ReLara’s advantage lies in its integration of the policy agent’s action into the reward construction, as evidenced in the RobotReach task. The target point for the robot is randomly selected in each episode, this can be considered as a higher dimension for SASR, because of the changing target. In this case, ReLara achieves better results, with its access to action information. This can be improved in SASR by incorporating the action as an additional feature in the state vector.
5.2 Effect of Self-Adaptive Success Rate Sampling
SASR introduces a novel self-adaptive mechanism that balances exploration and exploitation by maintaining the randomness of the shaped rewards. Figure 4 illustrates the generated rewards at five stages of learning in the AntStand task. Figure 4 (bottom) represents the learning curve, and Figure 4 (top) illustrates the distribution of the generated rewards over the "height of the ant" feature, one of the dimensions in the state vector.
Figure 4: Distribution of the shaped rewards over the height of the ant robot in the AntStand task, at different stages of the training process.
As learning progresses, the evolving shaped rewards exhibit two crucial attributes: the values transition from random to meaningful, and the variance decreases from uncertain to deterministic. In the early stages of training, the shaped rewards involve significant randomness due to higher uncertainty. Although these random signals contain limited information, they prompt the agent to take tiny optimization steps in diverse directions, effectively shifting the anchors of the policies to various regions. This strategy broadens the actions generated by SAC’s stochastic policy, encouraging the agent’s exploration and increasing the diversity of collected samples. From another perspective, the perturbations on the rewards can be seen as a relaxation of the assumption that only novel states should be rewarded, instead, randomly reward any state, which also shows its empirical effectiveness. In the later stages, the rewards are observed to converge to closely match the height of the robot, a meaningful and intuition-aligned metric, which enhances the agent’s exploitation of its policy. This self-adaptive mechanism for balancing exploration and exploitation improves SASR’s ability to gather positive samples earlier, thereby facilitating a faster convergence to the optimal policy.
5.3 Ablation Study
We conduct ablation studies to investigate five key components of SASR: (1) sampling from Beta distributions versus using the deterministic success rates; (2) the retention rate $\phi$, which controls the volume of historical information; (3) the bandwidth $h$ of the kernel density estimation; (4) the number of dimensions $M$ of RFF features; and (5) the weight factor $\lambda$, which adjusts the scales of the shaped rewards. We select six representative tasks to report the experimental results, and the quantitative data is provided in Appendix C.
Sampling from Beta distributions. We examine a variant of SASR that omits sampling from the Beta distribution and instead directly uses the success rate $N_{S}(s_{i})/(N_{S}(s_{i})+N_{F}(s_{i}))$. Comparative results are shown in Figure 5. In the early stages of learning, limited experience makes this success rate an inaccurate estimate of the true value. Using this overly confident fixed value can mislead the agent. Additionally, removing the process of sampling from the Beta distribution disrupts the agent’s exploration encouraged by random rewards, leading to slower convergence. The results highlight the importance of sampling from the Beta distribution for effective learning.
Figure 5: The impact of omitting Beta distribution sampling on the performance of SASR.
Figure 6: Learning performance of SASR on different retention rates $\phi$.
Retention rate $\phi$. The retention rate controls the counts’ estimation, subsequently impacting the Beta distribution and sampled rewards. We assess various retention rates in Figure 6. A high retention rate, such as $\phi=1$, preserving all samples, results in a densely populated and redundant experience pool, causing the Beta distribution to be prematurely overconfident, which can degrade performance. Conversely, a lower retention rate $\phi=0.01$ requires more iterations to accumulate sufficient samples, thereby slowing the convergence to meaningful rewards and affecting the agent’s convergence speed. The results suggest that carefully balancing the retention rate is crucial, and promising future work could involve adaptive adjustment of the retention rate during the learning process.
Bandwidth $h$ of Gaussian kernel. We evaluate different bandwidths $h$ of the Gaussian kernel for density estimation, as shown in Figure 7. Beyond fixed bandwidths, we also test a linearly decreasing bandwidth configuration ($h:0.5\rightarrow 0.1$), which represents progressively increasing confidence in KDE. Results indicate that an optimal bandwidth selection is crucial. A small bandwidth $h=0.01$ results in farther distances between samples, leading many to have zero estimated density; while a large bandwidth $h=1$ causes the samples indistinguishable due to the overly flat kernel function. Both scenarios lead to suboptimal performance. The decreasing bandwidth setting does not yield significant improvement and tends to reduce stability due to inconsistent density estimations.
Figure 7: Learning performance of SASR on different bandwidths $h$ of Gaussian kernel.
Figure 8: Learning performance of SASR on different RFF feature dimensions $M$.
Figure 9: Learning performance of SASR on different shaped reward weight factors $\lambda$.
RFF feature dimensions $M$. Learning performance on different RFF feature dimensions $M$ is shown in Figure 8. SASR shows relatively low sensitivity on $M$, provided it is large enough to capture the complexity of the states. Results indicate that values like $M=500,1000,2000$ all yield comparable performance levels. However, a significantly lower dimension, such as $M=50$, degrades performance, which is also suggested in the original RFF study (Rahimi and Recht, 2007).
Shaped reward weight factor $\lambda$. We evaluate the impact of different shaped reward scales on learning performance, as shown in Figure 9. We observe that SASR performs better with intermediate values like $\lambda=0.4,0.6,0.8$. At $\lambda=0.2$, the minimal shaped reward scale reduces state differentiation, leading to suboptimal performance. At $\lambda=1$, aligning the shaped reward scale with the environmental reward introduces excessive uncertainty and randomness, which may overwhelm the feedback disturbances and hinder learning. The findings highlight that maintaining a balanced reward scale is important for achieving optimal learning outcomes.
6 Conclusion and Discussion
We propose SASR in this paper, a self-adaptive reward shaping algorithm based on success rates, to address the sparse-reward challenge. SASR achieves an exploration and exploitation trade-off mechanism by generating shaped rewards from evolving Beta distributions. Experiments demonstrate that this adaptability significantly enhances the agent’s convergence speed. Additionally, the implementation of KDE and RFF formulates a highly efficient, non-parametric, and learning-free approach to deriving Beta distributions. This mechanism also provides a sound alternative to traditional count-based RS strategies, adapting effectively to continuous environments. Our evaluations confirm the superior performance of SASR in terms of sample efficiency and learning stability, and its robustness across various continuous-control tasks.
While SASR is designed for environments with sparse rewards, in settings where the environmental rewards are dense, the introduction and maintenance of additional shaped rewards could be unnecessary. On the other hand, SASR operates on the premise that sparse rewards indicate task completion to align the success rate notion. This correlation is generally straightforward, yet it may not intuitively apply in all contexts. Exploring how to extend SASR to more general sparse and even dense reward scenarios presents a promising direction for further research. Moreover, the derivation of Beta distributions heavily relies on the samples in the success and failure buffers. Currently, our method does not capture the relationships and varying importance of different states within the same trajectory, making the algorithm relatively sensitive to the retention rate, a crucial hyperparameter for achieving promising performance. Therefore, developing an adaptive retention rate or better mechanisms for maintaining the two buffers are crucial avenues for future enhancements.
Appendix A Derivation of Random Fourier Features
We incorporate Random Fourier Features (RFF) [Rahimi and Recht, 2007] to approximate the kernel functions in the KDE process for the SASR algorithm. Let the original state be $k$-dimensional, denoted as $\bm{s}\in\mathbb{R}^{k}$, and the kernel function be $k(\bm{s}_{i},\bm{s}_{j})$. RFF approximates the kernel function by projecting the input $k$-dimensional space into a $M$-dimensional feature space using a mapping function $z:\mathbb{R}^{k}\rightarrow\mathbb{R}^{M}$. The RFF-based kernel function is then defined as follows:
$k(\bm{s}_{i},\bm{s}_{j})\approx z(\bm{s}_{i})^{T}z(\bm{s}_{j}),$ (12)
We provide the derivation of the RFF approximation in this section.
First, we clarify that RFF primarily targets shift-invariant kernels, that satisfy $k(\bm{s}_{i},\bm{s}_{j})=k(\bm{s}_{i}-\bm{s}_{j})$. Common shift-invariant kernels include Gaussian kernels, Laplacian kernels, and Cauchy kernels. Given a shift-invariant kernel function $k(\Delta)$, we perform the inverse Fourier transform:
[…]
$=\mathbb{E}_{\bm{w}}\big{[}e^{\mathsf{i}\bm{w}^{T}(\bm{s}_{i}-\bm{s}_{j})}\big{]},$ (14)
where we can consider $\bm{w}\sim p(\bm{w})$ based on the Bochner’s theorem, and $p(\bm{w})$ is called the spectral distribution of kernel function. For the three types of shift-invariant kernels, the corresponding spectral distributions are listed in Table 2:
Table 2: Some shift-invariant kernels and their associated spectral distributions.
Kernel Kernel function, $k(\bm{s}_{i}-\bm{s}_{j})$ Spectral density, $p(\bm{w})$
Gaussian $\exp\left(-\frac{\|\bm{s}_{i}-\bm{s}_{j}\|_{2}^{2}}{2}\right)$ $(2\pi)^{-\frac{M}{2}}\exp\left(-\frac{\|\bm{w}\|_{2}^{2}}{2}\right)$
Laplacian $\exp\big{(}-\|\bm{s}_{i}-\bm{s}_{j}\|_{1}\big{)}$ $\prod_{m=1}^{M}\frac{1}{\pi(1+w_{d}^{2})}$
Cauchy $\prod_{i=1}^{k}\frac{2}{\pi\big{(}1+(\bm{s}_{i}-\bm{s}_{j})^{2}\big{)}}$ $\exp\big{(}-\|\bm{w}\|_{1}\big{)}$
Next, we perform the Euler’s formula transformation, which retains only the cosine term since we are dealing with real-valued functions, the kernel function can be further derived as:
$k(\bm{s}_{i},\bm{s}_{j})$ $=\mathbb{E}_{\bm{w}}\big{[}e^{\mathsf{i}\bm{w}^{T}(\bm{s}_{i}-\bm{s}_{j})}\big{]}$ (15)
$=\mathbb{E}_{\bm{w}}\big{[}\cos(\bm{w}^{T}(\bm{s}_{i}-\bm{s}_{j}))\big{]}$ (16)
$=\mathbb{E}_{\bm{w}}\big{[}\cos(\bm{w}^{T}(\bm{s}_{i}-\bm{s}_{j}))\big{]}+\mathbb{E}_{\bm{w}}\big{[}\mathbb{E}_{b}[\cos(\bm{w}^{T}(\bm{s}_{i}+\bm{s}_{j})+2b)]\big{]}$ (17)
$=\mathbb{E}_{\bm{w}}\big{[}\mathbb{E}_{b}[\cos(\bm{w}^{T}(\bm{s}_{i}-\bm{s}_{j}))+\cos(\bm{w}^{T}(\bm{s}_{i}+\bm{s}_{j})+2b)]\big{]}$ (18)
$=\mathbb{E}_{\bm{w}}\big{[}\mathbb{E}_{b}[\sqrt{2}\cos(\bm{w}^{T}\bm{s}_{i}+b)\sqrt{2}\cos(\bm{w}^{T}\bm{s}_{j}+b)]\big{]},$ (19)
where $b\sim\text{Uniform}(0,2\pi)$. Equation 17 holds since $\mathbb{E}_{b\sim\text{Uniform}(0,2\pi)}\big{[}\cos(t+2b)\big{]}=0$ for any $t$. Equation 19 is obtained from $\cos(A-B)+\cos(A+B)=2\cos(A)\cos(B)$, where $A=\bm{w}^{T}\bm{s}_{i}+b$, $B=\bm{w}^{T}\bm{s}_{j}+b$.
We define the mapping $z_{\bm{w},b}(\bm{s})=\sqrt{2}\cos(\bm{w}^{T}\bm{s}+b)$, then the kernel function can be approximated by the inner product of two vectors and the expectation can be approximated by Monte Carlo sampling:
$k(\bm{s}_{i},\bm{s}_{j})$ $=\mathbb{E}_{\bm{w}}\big{[}\mathbb{E}_{b}[z_{\bm{w},b}(\bm{s}_{i})z_{\bm{w},b}(\bm{s}_{j})]\big{]}$ (20)
$\approx\frac{1}{M}\sum_{m=1}^{M}z_{\bm{w}_{d},b_{d}}(\bm{s}_{i})z_{\bm{w}_{d},b_{d}}(\bm{s}_{j})$ (21)
$=z(\bm{s}_{i})^{T}z(\bm{s}_{j}).$ (22)
Therefore, we have derived the mapping function $z(\bm{s})=\sqrt{2/M}\cos(\bm{W}^{T}\bm{s}+\bm{b})$, where $\bm{W}\in\mathbb{R}^{M\times k}$ and $\bm{b}\in\mathbb{R}^{M}$. The RFF-based kernel function can be approximated by the inner product of the mapped features in the $M$-dimensional space.
Appendix B Derivation of Computation Complexity
In this section, we derive the computational complexity to retrieve the success or failure counts $N_{S}$ and $N_{F}$ for each iteration. Suppose the buffer size of $\mathcal{D}_{X}$ is $D$, the batch size of $\mathcal{B}$ is $B$, the corresponding counts are retrieved by calculating:
[…]
where the mapping function is defined as:
$z(\bm{s})=\sqrt{\frac{2}{M}}\cos(\bm{W}^{T}\bm{s}+\bm{b}),\quad\bm{W}\in\mathbb{R}^{k\times M},\quad\bm{b}\in\mathbb{R}^{M}.$ (24)
For each state, the mapping function calculation involves:
  • Matrix multiplication $\bm{W}^{T}\bm{s}$: $kM$.
  • Addition $\bm{W}^{T}\bm{s}+\bm{b}$: $M$.
  • Cosine calculation $\cos(\bm{W}^{T}\bm{s}+\bm{b})$: $M$.
Therefore, the computational complexity for calculating $z(\bm{s})$ for one state is $O(kM)$.
For each pair of states $(\bm{s}_{i},\bm{s}_{j})$, calculating the kernel involves $M$ multiplications and $M-1$ additions, thus, the complexity is $O(M)$.
For each iteration, we calculate the RFF mapping for all states in the buffer and the batch and then compute the kernel between them. The complexities involve three parts: RFF mapping for the buffer, RFF mapping for the batch and kernel calculation:
$O(DkM)+O(BkM)+O(MDB).$ (25)
Since the first two terms $O(DkM)$ and $O(BkM)$ are dominated by the last term $O(MDB)$ when the buffer size and the batch size are large, the overall computational complexity to retrieve the corresponding counts can be approximated as $O(MDB)$.
Appendix C Supplementary Experimental Results for Ablation Study
We provide the detailed quantitative results of the ablation study in this section.
Table 3: Ablation study #1: The average episodic returns and standard errors of SASR and the variant without sampling from Beta distributions.
Tasks SASR (with sampling) SASR (without sampling)
AntStand $\bm{39.12\pm 2.86}$ $23.25\pm 2.12$
AntFar $\bm{73.92\pm 4.97}$ $37.86\pm 5.16$
HumanStand $\bm{42.63\pm 2.17}$ $9.30\pm 0.083$
HumanKeep $\bm{180.98\pm 4.40}$ $170.51\pm 10.68$
RobotReach $\bm{81.29\pm 6.52}$ $33.76\pm 4.41$
RobotPush $\bm{137.06\pm 12.66}$ $70.37\pm 11.97$
Table 4: Ablation study #2: The average episodic returns and standard errors of SASR with different retention rates.
Tasks $\phi=1$ $\phi=0.1$ (default) $\phi=0.01$
AntStand $17.98\pm 1.86$ $\bm{39.12\pm 2.86}$ $20.27\pm 2.08$
AntFar $19.95\pm 3.00$ $\bm{73.92\pm 4.97}$ $48.47\pm 5.28$
HumanStand $9.25\pm 0.08$ $\bm{42.63\pm 2.17}$ $19.17\pm 1.52$
HumanKeep $179.10\pm 7.06$ $\bm{180.98\pm 4.40}$ $180.13\pm 6.27$
RobotReach $54.07\pm 8.39$ $\bm{81.29\pm 6.52}$ $35.18\pm 5.95$
RobotPush $3.91\pm 1.22$ $\bm{137.06\pm 12.66}$ $1.88\pm 0.47$
Table 5: Ablation study #3: The average episodic returns and standard errors of SASR with different bandwidths $h$ of Gaussian kernel.
Tasks $h=0.01$ $h=0.1$ $h=0.2$ (default) $h=1$ $h=0.5\rightarrow 0.1$
AntStand $7.94\pm 0.28$ $16.77\pm 1.18$ $\bm{39.12\pm 2.86}$ $12.76\pm 0.40$ $19.00\pm 1.63$
AntFar $11.68\pm 0.96$ $44.31\pm 3.34$ $\bm{73.92\pm 4.97}$ $27.96\pm 2.46$ $72.71\pm 4.57$
HumanStand $9.31\pm 0.06$ $\bm{43.64\pm 1.94}$ $42.63\pm 2.17$ $9.32\pm 0.06$ $13.81\pm 0.93$
HumanKeep $178.51\pm 4.99$ $\bm{181.44\pm 4.73}$ $180.98\pm 4.40$ $178.22\pm 4.97$ $178.04\pm 5.12$
RobotReach $37.03\pm 4.19$ $72.30\pm 5.67$ $\bm{81.29\pm 6.52}$ $35.87\pm 4.57$ $49.47\pm 5.1$
RobotPush $0.40\pm 0.13$ $8.80\pm 4.44$ $\bm{137.06\pm 12.66}$ $0.12\pm 0.06$ $0.34\pm 0.13$
Table 6: Ablation study #4: The average episodic returns and standard errors of SASR with different RFF feature dimensions $M$.
Tasks $M=50$ $M=500$ $M=1000$ (default) $M=2000$
AntStand $3.56\pm 0.12$ $12.08\pm 1.36$ $39.12\pm 2.86$ $\bm{42.52\pm 3.14}$
AntFar $27.76\pm 3.24$ $36.31\pm 2.69$ $\bm{73.92\pm 4.97}$ $61.79\pm 4.63$
HumanStand $9.46\pm 0.05$ $25.42\pm 1.83$ $42.63\pm 2.17$ $\bm{52.70\pm 2.13}$
HumanKeep $178.51\pm 4.84$ $181.81\pm 4.60$ $\bm{180.98\pm 4.40}$ $180.58\pm 4.63$
RobotReach $65.39\pm 6.22$ $38.25\pm 4.37$ $\bm{81.29\pm 6.52}$ $45.98\pm 4.90$
RobotPush $33.58\pm 4.86$ $54.56\pm 6.66$ $\bm{137.06\pm 12.66}$ $45.15\pm 6.15$
Table 7: Ablation study #5: The average episodic returns and standard errors of SASR with different shaped reward weight factors.
Tasks $\lambda=0.2$ $\lambda=0.4$ $\lambda=0.6$ (default) $\lambda=0.8$ $\lambda=1.0$
AntStand $7.60\pm 0.73$ $17.25\pm 1.58$ $\bm{39.12\pm 2.86}$ $28.43\pm 2.07$ $2.82\pm 0.09$
AntFar $22.51\pm 2.60$ $34.04\pm 3.87$ $\bm{73.92\pm 4.97}$ $67.62\pm 4.89$ $18.05\pm 1.79$
HumanStand $9.53\pm 0.04$ $33.95\pm 1.90$ $\bm{42.63\pm 2.17}$ $30.65\pm 1.65$ $12.23\pm 0.42$
HumanKeep $\bm{182.57\pm 4.32}$ $180.22\pm 4.60$ $180.98\pm 4.40$ $179.74\pm 4.53$ $179.62\pm 4.70$
RobotReach $41.05\pm 4.17$ $\bm{82.33\pm 6.37}$ $81.29\pm 6.52$ $71.30\pm 6.54$ $15.89\pm 2.06$
RobotPush $9.18\pm 1.99$ $118.92\pm 10.79$ $\bm{137.06\pm 12.66}$ $65.51\pm 12.71$ $47.98\pm 9.64$
Appendix D Network Structures and Hyperparameters
D.1 Network Structures
Figure 10 illustrates the structures of the policy network and Q-network employed in our experiments. The agent utilizes simple multilayer perceptron (MLP) models for these networks. Given the use of stochastic policies, the policy network features separate heads to generate the means and standard deviations of the inferred normal distributions, which are then used to sample actions accordingly.
Figure 10: The structures of policy network and Q-network in our implementation.
D.2 Hyperparameters
We have observed that SASR demonstrated high robustness and was not sensitive to hyperparameter choices. Table 8 shows the set of hyperparameters that we used in all of our experiments.
Table 8: The hyperparameters used in the SASR algorithm.
Hyperparameters Values
reward weight $\lambda$ (default) 0.6
kernel function bandwidth 0.2
random Fourier features dimension $M$ 1000
retention rate $\phi$ (default) 0.1
discounted factor $\gamma$ 0.99
replay buffer size $|\mathcal{D}|$ $1\times 10^{6}$
batch size 256
actor module learning rate $3\times 10^{-4}$
critic module learning rate $1\times 10^{-3}$
SAC entropy term factor $\alpha$ learning rate $1\times 10^{-4}$
policy networks update frequency (steps) $2$
target networks update frequency (steps) $1$
target networks soft update weight $\tau$ $5\times 10^{-3}$
burn-in steps $5000$
Appendix E Compute Resources
The experiments in this paper were conducted on a computing cluster, with the detailed hardware configurations listed in Table 9. The computing time for the SASR algorithm in each task (running 1,000,000 steps) was approximately $6\pm 2$ hours.
Table 9: The compute resources used in the experiments
Component Specification
Operating System (OS) Ubuntu 20.04
Central Processing Unit (CPU) 2x Intel Xeon Gold 6326
Random Access Memory (RAM) 256GB
Graphics Processing Unit (GPU) 1x NVIDIA A100 20GB
Brand Supermicro 2022